A balloon is filled with helium to a volume of 183 L on the ground, where the pressure is 745mmHg

      and the temperature is 20.0ºC. When the balloon ascends to a height of 2 miles, where the pressure is only

      555mmHg and the temperature is 5.0ºC. What is the volume of the helium in the balloon in L? 


1
Expert's answer
2021-03-25T04:06:47-0400

P1V1/T1 = P2V2/T2

P1 = 745mmHg

V1 = 183 L

T1 = 20.0° C

P2 = 555mmHg

T2 = 5.0°C

= (745×183)/20 = 555V2/5

= 6816.75 = 111V2

V2 = 6816.75/111

= 61.41 L